) Find the dual program of the following linear programming problem.

π‘€π‘–π‘›π‘–π‘šπ‘–π‘§π‘’ 𝑃 = 16π‘₯ βˆ’ 2𝑦 βˆ’ 5𝑧

𝑠𝑒𝑏𝑗𝑒𝑐𝑑 π‘‘π‘œ

π‘₯ + 4𝑦 βˆ’ 𝑧 β‰₯ 120

π‘₯ + 𝑦 + 3𝑧 ≀ 130

π‘₯ β‰₯ 0, 𝑦 β‰₯ 0, 𝑧 𝑖𝑠 π‘’π‘›π‘Ÿπ‘’π‘ π‘‘π‘Ÿπ‘–π‘π‘‘π‘’π‘‘.


1
Expert's answer
2021-09-28T01:16:40-0400

All constraints can be converted toΒ "\geq" Β by multiplying by -1. So we have;

"\text{Min} ~~p=16x-2y-5z\\
\text{subject to}\\
~~~~~x+4y-z\geq120\\
~-x-~~y+3z\geq-130\\
\text{and } x,y\geq 0; z \text{is unrestricted}."

Since the primal has three variables and two constraints, then the dual will have two variables and three constraints. Also, theΒ "z"Β variable is unrestricted in the primal, therefore the third constraint in the dual shall be equality.

Dual program is;

"\text{Max } ~~p=1220a-130b\\
\text{subject to }\\
~~~~~-a-b~~~\leq16\\
~~~~~~~4a-b~~~\leq-2\\
~~~~~-a-3b~=-5\\
\text{and } a,b\geq0."